OpenStudy (leah8343):

5^-8 • 5^−4 (this is exponents) A. 5^32 B. 5^4 C. 5^−4 D. 5^−12

OpenStudy (anonymous):

are you solving it completely? and are there options?

OpenStudy (leah8343):

they're options I can choose from

OpenStudy (anonymous):

what are they?

Nnesha (nnesha):

when we multiply same base we should add their exponents\[\huge\rm x^m \times x^n=x^{m+n}\]

Nnesha (nnesha):

so \[5^{-8} \times 5^{-4} = ?\]